1/2

Heinrich Heine,  Heinrich Heine,
The Lorelei  Lurleia
tradukita de L. W. Garnham  tradukita de N. N. 01

I do not know what it signifies. That I am so sorrowful? A fable of old Times so terrifies, Leaves my heart so thoughtful.

Ignoro, quid id sibi velit, Tristissimus cur sim, Antiqui aevi fabellam Cur saepe volverim.

The air is cool and it darkens, And calmly flows the Rhine; The summit of the mountain hearkens In evening sunshine line.

Vesperascit et frigescit, Et Rhenus leniter it, Cacumen montis lucescit, Dum Phoebus occidit.

The most beautiful Maiden entrances Above wonderfully there, Her beautiful golden attire glances, She combs her golden hair.

Sedet in summo montis Virgo pulcherrima, Auro nitet gemma frontis, Se pectit auricoma.

With golden comb so lustrous, And thereby a song sings, It has a tone so wondrous, That powerful melody rings.

Aureolo pectine pectit, Carmen canens procul, Mirandum id habet modum Nec non virilem simul.

The shipper in the little ship It effects with woe sad might; He does not see the rocky slip, He only regards dreaded height.

In cymba navitam mille Angores feri tenent, Non videt scopulos ille, Ocli non si sursum vident.

…  …

2/2

I believe the turbulent waves Swallow the last shipper and boat; She with her singing craves All to visit her magic moat.

Opinor undas devorare Nautam cum navicula, Effecit solo canendo Lurleia id dea.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de L. W. Garnham.

Arg-2-14 (2003-10-11 23:04:57)

L.W. Garnham, Bachelor of Arts, LEGENDS OF THE RHINE; mentioned in: A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Latinan de N. N. 01.

Arg-2-179 (2010-02-11 14:00:52)

http://www.poezio.net/version?poem-id=2&version-id=,14,179